HUBBLE SPACE TELESCOPE DAILY REPORT #1339 PERIOD COVERED: 0000Z (UTC) 03/02/95 - 0000Z (UTC) 03/03/95 Daily Status Report as of 062/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/2 5094 (Composition of Gas in Individual Interstellar Clouds: Cycle 4 Observations) The High Resolution Spectrograph (Side 2) was used to look for interstellar absorption lines in the star Eta Lup (aka, HD143118) to determine the column densities of a number of atomic species in the cloud. The observations were executed as scheduled. No other anomalies were noted. 1.2 Completed FOS/BL 5658 (White Dwarf Calibration) The Faint Object Spectrograph (Blue Side) was used to make absolute sensitivity calibrations of the detectors using standard white dwarf star GD153. HSTAR #5140 reported that the FOS encountered a speed check failure at 061/0815Z. The status buffer error message indicated that the FOC was overloaded. No other anomalies were noted. 1.3 Completed WF/PC-2 6140 (WF/PC-2 Cycle 4 Calibration: Ramp filters, Internal Calibration) The WF/PC-2 was used to make Earth calibration images as part of a ramp filter calibration test. The observations were completed as planned and no problems were reported. 1.4 Completed WF/PC-2 6142 (WF/PC-2 Cycle 4 Calibration: Earth Flats, Smaller Set Part II) The WF/PC-2 was used to make Earth Flat calibration images for the ultraviolet and narrow-band filters to correct for the telescope's illumination pattern. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Two Sets of WF/PC-2 5369 (HST Medium-Deep Survey: Cycle 4 Medium, High Galactic Latitude) The WF/PC-2 made a series of observations of high latitude objects as part of a medium deep sky survey. The observations were taken in parallel with other prime science observations. The observations were completed as scheduled, and no anomalies were noted. 1.6 Completed FOS/BL 5460 (Constraints on Massive Star Evolution from Carbon and Oxygen Abundance) The FOS (Blue Side) was used to observe BR93 to obtain UV spectra of several Wolf-Rayet stars in the LMC to determine the C/He and O/C ratios. HSTAR #5143 reported that the FOS encountered a binary acquisition failure at 061/1451Z during observation #25. A routine status buffer error message was received. No other anomalies were noted. 1.7 Completed FOS/BL 5444 (A UV Spectral Library of Metal- Poor Massive Stars) The FOS (Blue Side) was used for a snapshot program to obtain ultraviolet spectra of NS172-66 a metal-poor star in the LMC. It will be useful for stellar atmosphere and stellar evolution studies. The observations completed without incident. 1.8 Completed FOC/96 5409 (Direct UV and Spectral Tomogrpah Imaging of Alpha Orionis, Cycle 4 High) The Faint Object Camera (f/96) was used to make ultraviolet observations of the super giant star Alpha Orionis (aka, HD39801, Betelgeuse). The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2 5371 (HST Medium Deep Survey: Cycle 4 Medium Low Galactic Latitude) The WF/PC-2 was used in parallel with the prime science observations to make low galactic latitude deep multicolor images of objects such as galaxies, and stars in our galaxy. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 7 Successful : 7 2.3 Significant HSTARs: None 2.4 Operations Notes: During the day, COSTAR was transitioned from hold to the FOC sensor mode. The data inversions during STGT operations continue. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-